Car A and car B are traveling at different speeds. Car A can travel 175 miles in the same time it takes car B to travel 140 miles. If car A travels 10 mph faster than car B, how fast (in mph) is car B traveling?

Answer by mananth(16949)

Car B -----------x km/h------------ distance = 140 km
car A (x+10) km/h------------------distance = 175 km
time is same for both cars
t=d/r
140/x = 175/(x+10)
140(x+10)=175x
140x+1400= 175x
35x=1400
x= 1400/35
x= 40 km/h => speed of Car B